OpenStudy (dtan5457):

Guide me through this trig problem. cos(x+pi/6)-1=cos(x-pi/6)

OpenStudy (dtan5457):

I'm assuming you solve for cos(x+pi/6) and cos(x-pi/6) first?

OpenStudy (dtan5457):

i'm getting cos(x)(sqrt3/2)-sin(x)(1/2)-1=cos(x)(sqrt3/2)-sin(x)(1/2)

OpenStudy (dtan5457):

so is this just....-1?

OpenStudy (dtan5457):

@Michele_Laino @TheSmartOne

OpenStudy (anonymous):

Check the right hand side again. \[\cos(x-y)=\cos x\cos y\color{red}+\sin x\sin y\]

OpenStudy (dtan5457):

oh yeah, right

OpenStudy (dtan5457):

so we can cross out cos(x) and (sqrt3/2)?

OpenStudy (anonymous):

Yes, your sine terms remain.

OpenStudy (dtan5457):

the (1/2) is crossed out as well?

OpenStudy (michele_laino):

the right side is: cos(x)(sqrt3/2) + sin(x)(1/2)

OpenStudy (anonymous):

No I meant the sine terms along with their coefficients. You have \[-\frac{1}{2}\sin x-1=\frac{1}{2}\sin x\] which you could divide through by \(\dfrac{1}{2}\) to get \[-\sin x-2=\sin x\]

OpenStudy (dtan5457):

so sin(x)=-1 which would be 3pi/2?

OpenStudy (anonymous):

That's one solution, yes. If you're restricted to \(0\le x<2\pi\), then you're done.

OpenStudy (dtan5457):

think i got it now, thank you.

OpenStudy (anonymous):

yw